ADVANCED DYNAMICS — PHY 4241/5227 SOLUTIONS – SET 8

(24) SI Units. For up-to-date information visit the NIST website.

1. The second [s] is defined, so that the frequency of light between the two hyperfine levels of the ground state of the cesium ^{132}Cs atom is exactly 9,192,631,770 cycles per second.

2. The meter [m[ is defined, so that the speed of light in vacuum is exactly 299,792,458 [m/s].

3. A one kilogram weight is kept at Sèvres near Paris.

4. 1 N = 1 kg · m/s^{2}.

5. If two very long parallel wires one meter apart carry equal currents, the current in each is defined to be one Ampere [A] when the force per unit length on each wire is 2 × 10^{−7} [N/m].

6. One Coulomb [C] is the charge flowing through a cross-sectional area of a wire in one second when the current in the wire is one Ampere: C = A · s.

7. μ_{0} = 4π × 10^{−7} N/A^{2}.

8. ϵ_{0} = 8.85418782 × 10^{−12} C^{2}/(N · m^{2}).